# Patrick Cheridito and Moritz Weiss

### v1

@@ Abstract
In this paper, we introduce a novel reinforcement learning framework for optimal trade execution in a limit order book. We formulate the trade execution problem as a dynamic allocation task whose objective is the optimal placement of market and limit orders to maximize expected revenue. By employing multivariate logistic-normal distributions to model random allocations, the framework enables efficient training of the reinforcement learning algorithm. Numerical experiments show that the proposed method outperforms traditional benchmark strategies in simulated limit order book environments featuring noise traders submitting random orders, tactical traders responding to order book imbalances, and a strategic trader seeking to acquire or liquidate an asset position.
@@ 1 Introduction
The trade execution problem is a key concern for traders attempting to minimize trading costs and maximize execution speed. Its mathematical formulation goes back to Bertsimas and Lo (1998) and Almgren and Chriss (2001). Subsequent works such as Almgren (2003, 2012); Gatheral and Schied (2011, 2013); Obizhaeva and Wang (2013); Cheridito and Sepin (2014); Guéant and Lehalle (2015); Cartea and Jaimungal (2016); Curato et al. (2017); Di Giacinto et al. (2024) have generalized the price dynamics and form of the trade impact. But they all work with a simplified model of asset prices and market impact to keep the optimization problem tractable. A notable exception is the work Cartea and Jaimungal (2015), which studies an optimal execution problem with limit and market orders.
In this paper, we model the full limit order book and allow the strategy to submit market orders, limit orders, and cancellations. Since this results in high-dimensional state and action spaces, we train a reinforcement learning algorithm to find an optimal execution strategy. Various authors have applied reinforcement learning to trade execution using market simulations based on historical data. But typically, they impose certain restrictions on how orders can be placed. For instance, in Nevmyvaka et al. (2006), the entire inventory is replaced in every step, ignoring the queue positions [^1] Queue positions will be explained in more detail below. of limit orders. Moallemi and Wang (2022) frames the trade execution problem as an optimal stopping problem, determining the optimal time to execute a market order. Other approaches, such as Hendricks and Wilcox (2014); Ning et al. (2021); Leal et al. (2022) develop optimal execution schedules, but are not concerned with order placement.
Modeling market impact in simulations based on historical data is inherently limited, as such data is static and cannot react to artificial trades. While direct market impact, from consuming visible liquidity, can be approximated using limit order book data, indirect impact, which results from market participants adjusting to an algorithm’s behavior, cannot. As a result, many reinforcement learning approaches model only direct impact or ignore market impact altogether. A more realistic alternative uses stochastic simulations with interacting trading agents, where both direct and indirect impacts emerge naturally. Typically, these agents trade according to Poisson processes as in Cont et al. (2010). Other works use hand-crafted trading agents such as Byrd et al. (2020). Algorithms trained in such simulations (Hafsi and Vittori, 2024; Byun et al., 2023) still face limitations due to restricted state and action spaces.
We address the issues highlighted above. The main contributions of our paper are:
@@
  • A detailed empirical analysis of the algorithm’s performance in simulated markets. The simulations contain trading agents reacting to the algorithm’s order placement, causing direct and indirect market impact. The performance of the algorithm is measured against common trade execution benchmarks used by practitioners.
@@ 2 Limit order books
A limit order book is a mechanism used by most trading exchanges to process orders submitted by market participants. Market participants may submit market orders, limit orders, or cancellations. Arriving market orders are immediately matched with limit orders on the opposite side of the book. A limit order allows an agent to post a buy or sell offer behind the current best prices but poses the risk that the order is not filled. Cancellations allow an agent to withdraw active limit orders from the book.
Prices are given in discrete units called ticks [^2] A tick is the minimum price increment. It is typically 0.01 or 0.05 USD for common equity markets.. The best bid price $p^{b}(t)$ is the highest price at which buy orders are available in the limit order book. Conversely, the best ask price $p^{a}(t)$ is the lowest price at which sell orders are present in the book. One always has $p^{b}(t)<p^{a}(t).$ Usually, the mid-price $p(t)=(p^{b}(t)+p^{a}(t))/2$ is used as a reference price. The spread is given by $s(t)=p^{a}(t)-p^{b}(t),$ and is a common indicator of market liquidity. An order book can be described by a vector $v(t)=(v^{b}(t),v^{a}(t))$ that describes the volume on each price level, where
$v^{b}(t)=(v^{b,1}(t),\dots,v^{b,D}(t))\quad\mbox{and}\quad v^{a}(t)=(v^{a,1}(t),\dots,v^{a,D}(t)),$ (1)
for $D\in\operatorname{\mathbb{N}}.$ For $k\in\{1,\dots,D\}$, the quantity $v^{b,k}(t)$ is the volume at $k-1$ ticks below the best bid price. In particular, $v^{b,1}(t)$ is the volume at the best bid price. The volumes on the ask side $v^{a}(t)$ are defined analogously. The volume vector $v(t)$ reveals information about demand and supply in the market. For example, if there is more volume on the bid than the ask side, it is more likely that the mid-price will go up next. The volumes on each price level consist of individual limit orders by market participants and form a queue. If an incoming market order arrives at a price level, the limit orders that arrived first in the queue will be matched first. Any optimal execution strategy should keep track of price levels and queue positions of active limit orders. The lower the price level and the queue position of a limit order, the more likely it will be filled. For an extensive overview of limit order books we refer to the review paper Gould et al. (2013) or the textbook Bouchaud et al. (2018).
@@ 3 The trade execution problem
We consider a trade execution algorithm selling $M\in\operatorname{\mathbb{N}}$ lots [^3] A lot refers to a single unit of a financial asset. over a given time interval $[0,T]$ at discrete time steps $t_{n}=n\Delta t$ for $n\in\{0,1,\dots,N-1\},$ where $N\in\operatorname{\mathbb{N}},$ and $\Delta t=T/N.$ At each time $t_{n},$ the algorithm observes a market state $s_{n},$ takes an action $a_{n},$ and receives a random reward $r_{n}(s_{n},a_{n}).$ We assume at each time $t_{n},$ the algorithm takes into account what happened during the preceding subinterval $(t_{n-1},t_{n}],$ and takes an action $a_{n}$ immediately after $t_{n}.$ The states reflect the condition of the order book, including the current time, while the actions describe the algorithm’s order placement or cancellation. The algorithm learns an optimal policy, which we model as a conditional density $\pi$, such that
$J(\pi)=\underset{a\sim\pi}{\operatorname{\mathbb{E}}}\left[\sum_{n={0}}^{N-1}r(s_{n},a_{n})\right],$ (2)
is maximized. The subscript $a\sim\pi$ means that each action $a_{n}$ in the sequence of actions $a=(a_{0},a_{1},\dots,a_{N-1})$ is sampled from a conditional distribution with density $\pi(\cdot\mid s_{n})$ in each step $n\in\{0,1,\dots,N-1\}.$ The transition from one state $s_{n}$ to the next $s_{n+1}$ results from the interaction of our algorithm with the other market participants.
@@ 3.1 State space
  • The first $K-1$ entries of both volume vectors defined in (1), that is $(v^{b,1}(t),\dots,v^{b,K-1}(t))$ and $(v^{a,1}(t),\dots,v^{a,K-1}(t))$ for $K\in\operatorname{\mathbb{N}}$ with $K\leq D.$
  • The market order flow $\Delta m(t),$ which is the difference between the total volume of market buy and sell orders in the time interval $(t-\Delta t,t].$
  • The limit order flow $\Delta l(t),$ which is the difference between the total volume of limit buy and sell orders in the time interval $(t-\Delta t,t].$
  • The mid-price drift from time $t-\Delta t$ to $t,$ given by $\Delta p(t)=p(t)-p(t-\Delta t).$
@@
  • The time $t$ within the execution horizon.
  • The inventory $M(t),$ which is the number of lots the algorithm still has to sell.
  • The number of lots $m(t)\in\{0,1,\dots,M(t)\},$ which are currently sitting in the order book.
  • For $k\in\{1,2,\dots,m(t)\},$ the levels and queue positions $\left(l^{k}(t),q^{k}(t)\right)\in\operatorname{\mathbb{N}}\times\operatorname{\mathbb{N}}$ of the active limit orders. If $\left(l^{k}(t),q^{k}(t)\right)=(l,q),$ this means that the $k$-th active limit order is located $l^{k}(t)-1$ ticks above the best ask price, and at position $q^{k}(t)$ in the order queue.
We normalize the features above to speed up the learning process. More detailed explanations are given in Appendix D below.
Figure 1 illustrates the state space. The best bid and ask are $p^{b}(t)=100$ and $p^{a}(t)=101.$ The volumes on the bid and ask side are $(v^{b,1}(t),v^{b,2}(t),v^{b,3}(t),v^{b,4}(t))=(3,4,6,5)$ and $(v^{a,1}(t),v^{a,2}(t),v^{a,3}(t),v^{a,4}(t))=(2,4,5,7).$ The algorithm has $m(t)=2$ lots placed in the order book. The first lot has level and queue position $(l^{1}(t),q^{1}(t))=(2,3),$ and the second lot has level and queue position $(l^{2}(t),q^{2}(t))=(3,5).$
Figure 1: Limit order book with the algorithm’s orders in orange.
@@ 3.2 Action space
$\operatorname{\mathbb{S}}^{K}=\left\{a=(a^{0},a^{1},\dots,a^{K}),\sum_{k=0}^{K}a^{k}=1,a^{k}\geq 0\right\}.$
An action $a=(a^{0},a^{1},\dots,a^{K}),$ represents the allocation of orders to limit and market orders. More precisely, $a^{0}$ is the fraction of the remaining inventory $M(t)$ to be sent as a market order, $a^{k}$ for $k\in\{1,2,\dots K-1\}$ is the fraction of the remaining inventory $M(t)$ to be placed $k$ ticks above the best bid price, and $a^{K}$ is the fraction of the remaining inventory $M(t)$ to be held outside of the order book. We always require that $K\leq D,$ such that the algorithm always posts orders within the visible order book levels defined in (1). For illustration, let us assume $K=3,$ that the current inventory is $M(t)=10,$ and that $a(t)=[0.1,0.5,0.3,0.1].$ With this allocation, the algorithm sends one lot as a market order, five lots as a limit order one tick above the best bid price, three lots as a limit order two ticks above the best bid price, and keeps one lot on the side without placing it in the order book. In the event that the vector $a(t)M(t)$ includes non-integer values, we round the entries $a^{0}(t)M(t),a^{1}(t)M(t),\dots,a^{K-1}(t)M(t)$ to the closest integer sequentially, always assuring that total allocated orders to not exceed $M(t).$ To preserve queue positions, the algorithm only cancels orders if necessary to reach the new allocation when reallocating orders, while prioritizing canceling orders with higher queue positions first.
@@ 3.3 Rewards
In each time step $t_{n},$ the algorithm places market and limit orders according to an action $a_{n}.$ The market orders are filled immediately after $t_{n},$ leading to an instant cash flow, while the limit orders remain in the book waiting to be filled. The total cash flow from market and limit order fills in the interval $(t_{n},t_{n+1}]$ is denoted by $\bar{r}(s_{n},a_{n}).$ If any inventory is left at time $t_{N},$ the algorithm is forced to send a market order of size $M(t_{N}).$ In that case, the cash flow $\bar{r}(s_{N-1},a_{N-1})$ also contains the proceeds generated by the market order sent at time $t_{N}.$ Suppose that the algorithm sells $\gamma_{n}$ lots in the interval $(t_{n},t_{n+1}].$ Then, we define a normalized reward function, which compares the reward $\bar{r}(s_{n},a_{n})$ to the cash flow that would have been generated by trading at the initial bid price. More precisely, we define
@@
$J(\pi)=\frac{1}{M}\underset{a\sim\pi}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}\bar{r}(s_{n},a_{n})\right]-p^{b}(0),$ (3)
@@
Rather than forcing the algorithm to sell any leftover inventory, one could also introduce a soft constraint by using a penalty for the leftover inventory. However, forcing liquidation at terminal time is common in academic papers (Almgren and Chriss, 2001; Nevmyvaka et al., 2006) and is also a reasonable assumption for real-world applications. It is fairly common for traders to set a terminal time at which a trade execution algorithm must complete. For example, depending on the size of the initial inventory, a trader might set the time limit for full liquidation to 10 minutes or the end of the trading day.

### v2

# Reinforcement Learning for Trade Execution

@@ Abstract
In this paper, we introduce a novel reinforcement learning framework for optimal trade execution in a limit order book. We formulate the trade execution problem as a dynamic allocation task whose objective is the optimal placement of market and limit orders to maximize expected revenue. By modeling market and limit order allocations with multivariate logistic-normal distributions, the framework enables efficient training of the reinforcement learning algorithm. Numerical experiments show that the proposed method outperforms traditional benchmark strategies in simulated limit order book environments featuring noise traders submitting random orders, tactical traders responding to order book imbalances, and a strategic trader seeking to acquire or liquidate an asset position.
Keywords: Trade execution; Limit order book; Market order; Limit order; Reinforcement learning
@@ 1 Introduction
The trade execution problem is a key concern for traders attempting to minimize trading costs and maximize execution speed. Its mathematical formulation goes back to Bertsimas and Lo (1998) and Almgren and Chriss (2001). Subsequent works such as Almgren (2003, 2012); Gatheral and Schied (2011, 2013); Obizhaeva and Wang (2013); Cheridito and Sepin (2014); Guéant and Lehalle (2015); Cartea and Jaimungal (2016); Curato et al. (2017); Di Giacinto et al. (2024) have generalized the price dynamics and form of the trade impact. These works all focus on finding an optimal execution schedule. Related works such as Guilbaud and Pham (2013, 2015); Cartea and Jaimungal (2013, 2015); Huitema (2014); Cartea et al. (2015) investigate the optimal placement of market and limit orders, but they model the mid-price together with a reduced-form market impact to keep the optimization problem tractable.
In contrast to these approaches, our work studies the optimal use of market and limit orders, cancellations and their interactions with the full limit order book. Since this results in high-dimensional state and action spaces, we train a reinforcement learning algorithm to find an optimal execution strategy. Various authors have applied reinforcement learning to trade execution using market simulations based on historical data. But typically, they impose certain restrictions on how orders can be placed. For instance, in Nevmyvaka et al. (2006), the entire inventory is replaced in every step, ignoring the queue positions of limit orders. Moallemi and Wang (2022) frame the trade execution problem as an optimal stopping problem, determining the optimal time to execute a market order. Other approaches, such as Hendricks and Wilcox (2014); Ning et al. (2021); Leal et al. (2022) develop optimal execution schedules, but are not concerned with order placement.
Modeling market impact in simulations based on historical data faces the limitation that the data is given and does not react to simulated trades. While direct market impact due to the consumption of liquidity on the other side of the book can be modelled using real limit order book data, indirect impact, which results from market participants adjusting to an algorithm’s behavior, cannot. As a result, many reinforcement learning approaches model only direct impact or ignore market impact altogether. A more realistic alternative uses stochastic simulations with interacting trading agents, where both direct and indirect impacts emerge naturally. Typically, these agents trade according to Poisson processes as in Cont et al. (2010). Other works use hand-crafted trading agents such as Byrd et al. (2020). Algorithms trained in such simulations (Hafsi and Vittori, 2024; Byun et al., 2023) still face limitations due to restricted state and action spaces.
We introduce a new approach. Our main contributions are:
@@
  • A detailed empirical analysis of the algorithm’s performance in simulated markets. The simulations contain trading agents reacting to the algorithm’s order placement, causing direct and indirect market impact. The algorithm’s performance is compared with common trade execution benchmarks used by practitioners.
@@ 2 Limit order books
A limit order book is a mechanism used by most trading exchanges to process orders submitted by market participants. Market participants may submit market orders, limit orders, or cancellations. Arriving market orders are immediately matched with limit orders on the opposite side of the book. Limit orders allow agents to post buy or sell offers away from the best market offers, but they are not guaranteed to be filled. Cancellations allow an agent to withdraw active limit orders from the book. Prices are given in discrete units called ticks [^1] A tick is the minimum price increment. It is typically 0.01 or 0.05 USD for common equity markets.. The best bid price $p^{b}(t)\in\operatorname{\mathbb{R}}_{+}$ is the highest price at which buy orders are available in the limit order book. Conversely, the best ask price $p^{a}(t)\in\operatorname{\mathbb{R}}_{+}$ is the lowest price at which sell orders are present in the book. One always has $p^{b}(t)<p^{a}(t).$ Usually, the mid-price $p(t)=(p^{b}(t)+p^{a}(t))/2$ is used as a reference price. The spread is given by $s(t)=p^{a}(t)-p^{b}(t)\in\operatorname{\mathbb{R}}_{+}$ and is a common indicator of market liquidity. An order book can be described by a vector $v(t)=(v^{b}(t),v^{a}(t))$ that describes the volume on each price level, where
$v^{b}(t)=(v^{b,1}(t),\dots,v^{b,D}(t))\in\operatorname{\mathbb{R}}^{D}_{+}\quad\mbox{and}\quad v^{a}(t)=(v^{a,1}(t),\dots,v^{a,D}(t))\in\operatorname{\mathbb{R}}^{D}_{+}$ (1)
for $D\in\operatorname{\mathbb{N}}.$ For $k\in\{1,\dots,D\}$, the quantity $v^{b,k}(t)$ is the volume k1k-1 ticks below the best bid price. In particular, $v^{b,1}(t)$ is the volume at the best bid price. The volumes on the ask side $v^{a}(t)$ are defined analogously. The volume vector $v(t)$ reveals information about demand and supply in the market. For example, if there is more volume on the bid than the ask side, it is more likely that the mid-price will go up next. The volumes on each price level consist of individual limit orders by market participants and form a queue. If an incoming market order arrives at a price level, the limit orders that arrived first in the queue will be matched first. Any optimal execution strategy should keep track of price levels and queue positions of active limit orders. For instance, the lower the price level and the queue position of a limit order, the more likely it will be filled. For an extensive overview of limit order books we refer to the review paper Gould et al. (2013) or the textbook Bouchaud et al. (2018).
@@ 3 The trade execution problem
We consider a trade execution algorithm selling $M\in\operatorname{\mathbb{N}}$ lots [^2] A lot refers to a single unit of a financial asset. over a given time interval $[0,T]$ at discrete time steps $t_{n}=n\Delta t$ for $n\in\{0,1,\dots,N-1\},$ where $N\in\operatorname{\mathbb{N}},$ and $\Delta t=T/N.$ We denote the algorithm’s inventory at time $t\in\{t_{0},t_{1},\dots,t_{N}\}$ by $M(t)\in\{0,1,\dots,M\}.$ The algorithm starts with an initial inventory $M(0)=M$ and must sell everything by the final time $T$, so that $M(T)=0$. At each time $t_{n},$ the algorithm observes a market state $s_{n},$ takes an action $a_{n},$ and receives a random reward $r_{n}(s_{n},a_{n}).$ We assume at each time $t_{n},$ the algorithm takes into account what happened during the preceding subinterval $(t_{n-1},t_{n}],$ and takes an action $a_{n}$ immediately after $t_{n}.$ The states reflect the condition of the order book, including the current time, while the actions describe the algorithm’s order placement or cancellation. The algorithm learns an optimal policy, which we model as a conditional density $\pi$ on the action space, such that
$J(\pi)=\underset{\begin{subarray}{c}a\sim\pi_{\theta}\\ s_{0}\sim\rho\end{subarray}}{\operatorname{\mathbb{E}}}\left[\sum_{n={0}}^{N-1}r(s_{n},a_{n})\right]$ (2)
is maximized. The subscript $a\sim\pi$ means that each action $a_{n}$ in the sequence of actions $a=(a_{0},a_{1},\dots,a_{N-1})$ is sampled from a conditional distribution with density $\pi(\cdot\mid s_{n})$ in each step $n\in\{0,1,\dots,N-1\}.$ The transition from one state $s_{n}$ to the next $s_{n+1}$ results from the interaction of our algorithm with the other market participants. The subscript $s_{0}\sim\rho$ means that the initial state $s_{0}$ is sampled from an initial distribution $\rho.$
@@ 3.1 State space
  • The first K1K-1 entries of both volume vectors defined in (1), that is $(v^{b,1}(t),\dots,v^{b,K-1}(t))$ and $(v^{a,1}(t),\dots,v^{a,K-1}(t))$ for $K\in\operatorname{\mathbb{N}}$ with $K-1\leq D.$
  • The market order flow $\Delta^{M}(t)\in\operatorname{\mathbb{R}},$ which is the difference between the total volume of market buy and sell orders in the time interval $(t-\Delta t,t].$
  • The limit order flow $\Delta^{L}(t)\in\operatorname{\mathbb{R}},$ which is the difference between the total volume of limit buy and sell orders in the time interval $(t-\Delta t,t].$
  • The cancellation order flow $\Delta^{C}(t)\in\operatorname{\mathbb{R}},$ which is the difference between the total volume of sell and buy limit order cancellations in the interval $(t-\Delta t,t]$.
  • The mid-price drift from time $t-\Delta t$ to $t,$ given by $\Delta p(t)=p(t)-p(t-\Delta t)\in\operatorname{\mathbb{R}}.$
@@
  • The time $t\in\{t_{0},t_{1},\dots,t_{N-1}\}$ within the execution period.
  • The inventory $M(t)\in\{1,2,\dots,M\},$ which is the number of lots the algorithm still has to sell.
  • The number of limit orders $m(t)\in\{0,1,\dots,M(t)\},$ currently resting in the limit order book.
  • For $i\in\{1,2,\dots,m(t)\},$ the levels and queue positions $\left(l^{i}(t),q^{i}(t)\right)\in\operatorname{\mathbb{N}}\times\operatorname{\mathbb{N}}$ of the active limit orders. E.g. $\left(l^{i}(t),q^{i}(t)\right)=(l,q)$ means that the $i$-th active limit order is located $l^{i}(t)-1$ ticks above the best ask price, and at position $q^{i}(t)$ in the order queue.
We normalize the features above to speed up the learning process. Furthermore, we add an additional feature that represents the allocation of the algorithm’s orders per price level. Both are explained in more detail in Appendix B.3 below.
@@ 3.2 Action space
$\operatorname{\mathbb{S}}^{K}=\left\{a=(a^{0},a^{1},\dots,a^{K}):\,\sum_{k=0}^{K}a^{k}=1,a^{k}\geq 0\right\}.$
An action $a=(a^{0},a^{1},\dots,a^{K}),$ represents the allocation of orders to market and limit orders. More precisely, $a^{0}$ is the fraction of the remaining inventory $M(t)$ to be sent as a market order, $a^{k}$ for $k\in\{1,2,\dots K-1\}$ is the fraction of the remaining inventory $M(t)$ to be placed $k$ ticks above the best bid price, and $a^{K}$ is the fraction of the remaining inventory $M(t)$ to be held outside of the order book. We require that $K-1\leq D,$ such that the algorithm always posts orders within the visible order book levels defined in (1). If the vector $a(t)M(t)$ includes non-integer values, we round the entries $a^{0}(t)M(t),a^{1}(t)M(t),\dots,a^{K-1}(t)M(t)$ to the closest integer sequentially, always assuring that total allocated orders to not exceed $M(t).$ To preserve queue positions, the algorithm only cancels orders if necessary to reach the new allocation when reallocating orders, prioritizing canceling orders with higher queue positions first.
Figure 1 illustrates the states and actions. At time $t,$ the entire inventory $M(t)=m(t)=5$ is sitting as limit orders in the limit order book. The algorithm then decides to reallocate the orders according to the action $a(t)=(0.4,0.2,0.2,0.2,0.0).$ The left panel of the figure shows the order book before the action. The right panel shows the order book at time $t+$, immediately after the action $a(t).$ The best bid and ask prices are $p^{b}(t)=100$ and $p^{a}(t)=101.$ The bid and ask volumes before taking the action are $(v^{b,1}(t),v^{b,2}(t),v^{b,3}(t))=(3,4,6)$ and $(v^{a,1}(t),v^{a,2}(t),v^{a,3}(t))=(2,5,6).$ The volumes right after the action are given by $(v^{b,1}(t+),v^{b,2}(t+),v^{b,3}(t+))=(1,4,6)$ and $(v^{a,1}(t+),v^{a,2}(t+),v^{a,3}(t+),v^{a,4}(t+))=(2,4,5).$ The algorithm’s orders are colored in orange. The buy and sell limit orders of other market participants are shown in blue and red, respectively. To reach the new allocation, the algorithm must cancel one order on the second and third ask levels. Each of those levels contains two orders. On each price level, the algorithm cancels the order with the worst queue position, as indicated by the two crossed orange boxes. Here, we assume that it is always better to cancel orders with a higher queue position first. After canceling the two limit orders, it sends a market sell order of size two, as indicated by the crossed blue box. The levels and queue position of the algorithm at time $t$ are given by $(l^{1}(t),q^{1}(t))=(1,2),(l^{2}(t),q^{2}(t))=(2,1),(l^{3}(t),q^{3}(t))=(2,3),(l^{4}(t),q^{4}(t))=(3,2),(l^{5}(t),q^{5}(t))=(3,5).$ The levels and queue positions at time $t+$ are given by $(l^{1}(t+),q^{1}(t+))=(1,2),(l^{2}(t+),q^{2}(t+))=(2,1),(l^{3}(t+),q^{3}(t+))=(3,2).$ The inventory and the number of limit orders resting in the book are $M(t+)=m(t+)=3.$
Figure 1: Limit order book with the algorithm’s orders in orange. Limit buy and sell orders are colored in blue and red, respectively. The algorithm takes action $a(t).$ The left panel shows the order book before the action. The right panel shows the order book after the action. The crosses indicate order cancellations (for the orange orders) or market order fills (for the blue orders).
@@ 3.3 Rewards
At each time $t_{n},$ the algorithm places market and limit orders according to an action $a_{n}.$ The market orders are filled immediately after $t_{n},$ leading to an instant cash flow, while the limit orders remain in the book waiting to be filled. The total cash flow from market and limit order fills in the interval $(t_{n},t_{n+1}]$ is denoted by $\bar{r}(s_{n},a_{n}).$ If any inventory is left at time $t_{N},$ the algorithm is forced to send a market order of size $M(t_{N}).$ In that case, the cash flow $\bar{r}(s_{N-1},a_{N-1})$ also contains the proceeds generated by the market order sent at time $t_{N}.$ Suppose that the algorithm sells $\gamma_{n}$ lots in the interval $(t_{n},t_{n+1}].$ Then, we define a normalized reward function, which compares the reward $\bar{r}(s_{n},a_{n})$ to the cash flow that would have been generated by trading at the initial bid price. More precisely, we define
@@
$J(\pi)=\frac{1}{M}\underset{\begin{subarray}{c}a\sim\pi_{\theta}\\ s_{0}\sim\rho\end{subarray}}{\operatorname{\mathbb{E}}}\left[\sum_{n=0}^{N-1}\bar{r}(s_{n},a_{n})\right]-p^{b}(0),$ (3)
@@
Rather than forcing the algorithm to sell any leftover inventory, one could also introduce a soft constraint by using a penalty for the leftover inventory. However, forcing liquidation at terminal time is common in academic papers (Almgren and Chriss, 2001; Nevmyvaka et al., 2006) and is also a reasonable assumption for real-world applications. It is fairly common for traders to set a time limit for a trade execution algorithm to complete.